OpenStudy (anonymous):

after a 30% reduction, you purchase a new soda on sale for 315. what was the original price of the sofa?

OpenStudy (anonymous):

315*(1-.3)=original price =220.5

OpenStudy (anonymous):

x - 30/100 = 315 x = 315 + 30/100 x = ?

OpenStudy (anonymous):

oops I think I did that wrong

OpenStudy (anonymous):

x*(1-.3)=315 x=450

OpenStudy (anonymous):

how did u get 450 cause i got 409.5

OpenStudy (anonymous):

blender is right

OpenStudy (anonymous):

x=315/.7 7 into 31 4 times 7*4=28 remainder: 3, drop the 5 7 into 35, 5 times

OpenStudy (anonymous):

i am confused

OpenStudy (anonymous):

About which part, setting up the equation, reducing the equation, or dividing 315 by .7?

OpenStudy (anonymous):

pretty much all of it

OpenStudy (anonymous):

Ok we can go through it slowly. First, we now that the product x was reduced by 30%. This means that originally we had some x and muliplied it by .3 to get the discount. Instead of doing this separately, we take 70% of the original x to find the discounted price. So that is .7x, which we know to be equal to 315. Then just divide .7 on each side (remember % is just a fancy way of saying that number per (divided by) cent (100).

OpenStudy (anonymous):

Does that make more sense?

OpenStudy (anonymous):

yeah now it does

OpenStudy (anonymous):

Cool, goodluck with your math final, I gotta go study myself. you can message me on aim (ooblender) if you have any urgent questions.

OpenStudy (anonymous):

i have yahoo sorry

OpenStudy (anonymous):

zblender89@yahoo